# On novel Hamiltonian description of the nonholonomic Suslov problem

### v1

@@ 4 Conclusion
Suppose we are given a vector field $X$. We want to represent it in the Hamiltonian form (2.2)
$X=PdH\,,$
where $H$ and $P$ are scalar and tensor invariants, i.e. some solutions of the invariance equation (1.3)
$\mathcal{L}_{X}H=0\qquad\mbox{and}\qquad\mathcal{L}_{X}P=0\,.$
In addition we have to impose the Jacobi condition $[\![P,P]\!]=0$ on invariant bivector $P$.
In this paper we solve these equations for the nonholonomic Suslov problem and find a few cubic Poisson brackets which allows us to represent a given vector field $X$ (LABEL:ep-eq,3.1 in the Hamiltonian form.
The corresponding Poisson bivectors $P_{a}$ and $P_{b}$ (2.8,2.9) are rank four tensor fields, whereas the Poisson bivectors $P_{c},P_{d}$ (3.2) and $P_{e}$ are rank two tensor fields. These bivectors define cubic Poisson brackets and, therefore, divergency of $X$ does not vanish, as for the standard Hamiltonian equations of motion on the symplectic manifold $\mathbb{R}^{2n}$.
We also find the corresponding globally defined Casimir functions and prove that rank four solutions define regular Poisson manifolds that allows us to introduce Darboux coordinates on the corresponding symplectic leaves. For the rank two solutions we have only two globally defined Casimir functions and construction of their symplectic leaves requires further investigation.
The study was carried out with the financial support of the Ministry of Science and Higher Education of the Russian Federation in the framework of a scientific project under agreement No. 075-15-2024-631.

### v2

@@ 4 Conclusion
Our aim is to consider a system of ordinary differential equations
$\dot{\mathrm{x}}_{k}=X_{k}(\mathrm{x}_{1},\ldots,\mathrm{x}_{N})\,.$
having the same structure as the Euler-Poisson equations for the free top, for the Suslov problem and for some other known systems of equations from the textbook [3]. For all these systems we can divide $N$ variables $\mathrm{x}=(\mathrm{x}_{1},\ldots,\mathrm{x}_{N})$ on two parts
$x=(x_{1},\ldots,x_{n})\qquad\mbox{and}\qquad y=(y_{1},\ldots,y_{m})\,,\qquad n+m=N$
and consider a system of equations with quadratic right-hand sides
$\dot{x}_{k}=\sum_{i,j=1}^{n}a_{k}^{ij}x_{i}x_{j}\qquad\mbox{and}\qquad\dot{y}_{\ell}=\sum_{i=1}^{n}\sum_{j=1}^{m}b_{\ell}^{ij}x_{i}y_{j}\,,\qquad a_{k}^{ij},b_{\ell}^{ij}\in\mathbb{R}\,.$ (4.9)
We also suppose that first subsystem of equations has n1n-1 independent first integrals $f_{1},\ldots,f_{n-1}$ depending on $x_{k}$, whereas second subsystem of equations has m2m-2 independent first integrals $g_{1},\ldots,g_{n-2}$ depending on $y_{\ell}$.
The intersection of the invariant levels of the first integrals $f_{1},\ldots,f_{n}$ determine solutions $x_{k}(t)$. When we substitute these solutions into the second subsystem, we obtain non-autonomous system of $m$ equations
$\dot{y}_{\ell}=\sum_{i=1}^{n}\sum_{j=1}^{m}b_{\ell}^{ij}x_{i}(t)y_{j}\,,\qquad\ell=1,\ldots,m,$
which has m2m-2 first integrals $g_{1},\ldots,g_{m-1}$. It allows us to study solutions of this counterpart of the Poisson equations in the framework of the Carathéodory theory.
We want to find an invariant Hamiltonian representation of differential equations (4.9). In this note we consider two partial cases of equations (4.9) with $N=5$ and $N=6$. A general case will be considered in subsequent publication.
Funding
The study was carried out with the financial support of the Ministry of Science and Higher Education of the Russian Federation in the framework of a scientific project under agreement No 075-15-2025-013 by St. Petersburg State University as part of the national project “Science and Universities” in 2025.